# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 4.3. Injective non-Archimedean Polish abelian groups
(2$\Rightarrow$(3) If $A$ is divisible, then $A\cong\mathrm{\mathrm{lim}}_{k}A_{k}$ where, for every $k\in\omega$, $A_{k}$ is countable, divisible, and torsion-free and $A_{k+1}\rightarrow A_{k}$ is surjective with divisible kernel. Since countable divisible groups are injective for countable abelian groups, it follows that $A$ is isomorphic to the product of countable divisible groups.
@@
    The category $\mathbf{proLiePAb}$ has enough injective objects and enough projective objects, and homological dimension $1$;
@@
(1) The same proof as Corollary 4.17 shows that $\mathbf{proLiePAb}$ has enough projective objects and homological dimension $1$. The same proof as Corollary 4.6 shows that $\mathbf{proLiePAb}$ has enough projective objects.
(2) We have that $\mathbf{proLiePAb}_{\mathbb{Z}}$ has enough projectives by Lemma 3.39 and Theorem 4.16. Furthermore, it has homological dimension $1$ by Lemma 3.41. Suppose that $A$ is a type $\mathbb{Z}$ Polish abelian group. Suppose that $A\rightarrow D\rightarrow B$ is a short exact sequence where $D$ and $B$ are type $\mathbb{Z}$ Polish abelian groups, and $D$ is injective. Since $\mathbf{proLiePAb}_{\mathbb{Z}}$ has homological dimension $1$, it follows that $B$ is injective. Then for $p\in\mathbb{Z}$ we have that $\mathrm{Hom}\left(\mathbb{Z}/p\mathbb{Z},B\right)=0$ and $\mathrm{Ext}\left(\mathbb{Z}/p\mathbb{Z},D\right)=0$. This implies that $\mathrm{Ext}\left(\mathbb{Z}/p\mathbb{Z},A\right)=0$ and $A$ is divisible. ∎
@@ 4.4. Ext as a group with a Polish cover
By Theorem 3.47, a pro-Lie Polish abelian group is essentially injective if and only if $G$ is isomorphic to $T\oplus V\oplus A$ where $T$ is a torus, $V$ is a vector group, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We let $\mathcal{D}$ be the collection of essentially injective pro-Lie Polish abelian groups. By Lemma 3.51, every pro-Lie Polish abelian group is isomorphic to a closed subgroup of an element of $\mathcal{D}$. We will now show that the class $\mathcal{D}$ is closed under taking quotients by closed subgroups. Towards this goal, we isolate a few lemmas concerning $\mathrm{Ext}$ of countable abelian groups, regarded as a group with a Polish cover.
@@
Suppose that there exists a short exact sequence of pro-Lie Polish abelian groups $N\rightarrow V\oplus B\rightarrow K$ where $V$ is a vector group, $B$ is non-Archimedean, and $K$ is compact and connected. Then $K$ is a torus group.
Proof.
Let $F$ be a free countable abelian group and $\rho:F^{\omega}\rightarrow N$ be a surjective continuous homomorphism. Since the induced homomorphism $\mathrm{Ext}\left(K,F^{\omega}\right)\rightarrow\mathrm{Ext}\left(K,B\right)$ is surjective, we have a commuting diagram
$\begin{array}[]{ccccc}N&\rightarrow&V\oplus B&\rightarrow&K\\ \uparrow&&\uparrow&&\uparrow\\ F^{\omega}&\rightarrow&L&\rightarrow&K\end{array}$
where the right-most arrow is the identity of $K$, and the diagram
$\begin{array}[]{ccc}N&\rightarrow&V\oplus B\\ \uparrow&&\uparrow\\ F^{\omega}&\rightarrow&L\end{array}$
is a push-out in the category of pro-Lie Polish abelian groups. Since $\left(V\oplus B\right)_{\mathbb{S}_{1}}=0$ and $\left(F^{\omega}\right)_{\mathbb{S}^{1}}=0$, this implies that $L_{\mathbb{S}^{1}}=0$.
Considering the isomorphism $\mathrm{Ext}\left(K,F^{\omega}\right)\cong\mathrm{Ext}\left(K,F\right)^{\omega}$, we have extensions $F\rightarrow L_{i}\rightarrow K$ for $i\in\mathbb{N}$ such that $L\cong\prod_{i\in\mathbb{N}}L_{i}$. For $i\in\mathbb{N}$, we have that $\left(L_{i}\right)_{\mathbb{S}_{1}}=0$ and $L_{i}$ is locally compact, as locally compact groups form a thick subcategory of $\mathbf{PAb}$ [25, Theorem 6.17].
Thus, we have a short exact sequence $F\rightarrow L_{1}\rightarrow K$ $L_{1}$ is a locally compact Polish abelian group and $\left(L_{1}\right)_{\mathbb{S}_{1}}=0$. We can write $L_{1}=W\oplus\Delta$ where $W$ is a vector group and $\Delta$ is a totally disconnected locally compact Polish abelian group. Let $U\subseteq\Delta$ be a compact open subgroup of $\Delta$, and let $\pi:W\oplus\Delta\rightarrow K$ be the quotient map. Since $\pi$ is an open mapping, we have that $\pi\left(W\oplus U\right)$ is a compact open subgroup of $K$. Since $K$ is connected, this implies that $\pi\left(W\oplus U\right)=K$. Thus, after replacing $F$ with $F\cap\left(W\oplus U\right)$ and $L_{1}$ with $W\oplus U$, we can assume that $\Delta=U$ is compact, and hence profinite.
For a locally compact Polish abelian group $G$, we let $G^{\vee}$ be its Pontryagin dual. Since $F$ is a free countable abelian group, $F^{\vee}$ is a torus group. Since $U$ is a profinite Polish abelian group, $U^{\vee}$ is a locally finite countable abelian group. Since $K$ is compact and connected, $K^{\vee}$ is countable and torsion-free. From the short exact sequence $F\rightarrow W\oplus\Delta\rightarrow K$ one obtains by Pontryagin duality a short exact sequence
$K^{\vee}\rightarrow W^{\vee}\oplus\Delta^{\vee}\rightarrow F^{\vee}$
This short exact sequence induces an exact sequence
$0=\mathrm{Hom}\left(F^{\vee},\mathbb{Z}\right)\rightarrow\mathrm{Ext}\left(K^{\vee},\mathbb{Z}\right)\rightarrow\mathrm{Ext}\left(W^{\vee}\oplus\Delta^{\vee},\mathbb{Z}\right)\rightarrow\mathrm{Ext}\left(F^{\vee},\mathbb{Z}\right)=0\text{.}$
Thus, we have that
$\mathrm{Ext}\left(K^{\vee},\mathbb{Z}\right)\cong\mathrm{Ext}\left(W^{\vee}\oplus\Delta^{\vee},\mathbb{Z}\right)\cong\mathrm{Ext}\left(\Delta^{\vee},\mathbb{Z}\right)$
is a Polish group by Lemma 4.19. Since $K^{\vee}$ is torsion-free, this implies that $K^{\vee}$ is free abelian by Lemma 4.19. Therefore, we have that $K$ is a torus group. ∎
Lemma 4.22.
@@
Let $H$ be a pro-Lie Polish group. Then we have that the connected component $H_{0}$ of zero in $H$ is $H_{\mathbb{S}^{1}}\oplus H_{\mathbb{R}}$, and $H/H_{0}$ is non-Archimedean. Thus, we have that $H=V\oplus A$ for some vector group $A$ and non-Archimedean abelian Polish group $A$ if and only if $H_{\mathbb{S}^{1}}=0$.
Suppose that $G$ is an essentially injective pro-Lie Polish abelian group and $N$ is a closed subgroup of $G$. Set $H:=G/N$ and let $\pi:G\rightarrow H$ be the quotient map. Since $G$ is essentially injective, we have that $G=T\oplus V\oplus A$ where $V$ is a vector group, $T$ is a torus, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We need to prove that $G/N$ is injective.
We have that $T/\left(T\cap N\right)$ is injective, being the quotient of a torus group by a closed subgroup. Furthermore, we have that $\pi^{-1}\left(\pi\left(N\right)\right)=N+T$ is closed in $G$, and hence $\pi\left(N\right)$ is closed is $V\oplus A$. Thus, we have a short exact sequence
@@
Since $N$ is a closed subgroup of $G=V\oplus A$, we have that $c_{0}\left(N\right)\subseteq V=c_{0}(G)$ and hence $c_{0}\left(N\right)=N_{\mathbb{R}}$ is a vector group and $N_{\mathbb{S}^{1}}=0$. Thus, we can write $N=N_{\mathbb{R}}\oplus\Xi$ where $\Xi$ is non-Archimedean. Since $N_{\mathbb{R}}\subseteq V$ is a vector group, we can assume without loss of generality, possibly after replacing $G$ with $G/N_{\mathbb{R}}$, that $N_{\mathbb{R}}=0$ and $N=\Xi$ is non-Archimedean.
We can write $H=W\oplus S$ where $W$ is a vector group and $S$ has no nonzero closed vector subgroups. By projectivity of $W$, we can write $G\cong G^{\prime}\oplus W$ where $N\subseteq G^{\prime}$ and $G^{\prime}/N\cong S$. Thus, without loss of generality, after replacing $G$ with $G/W$, we can assume that $H$ has no nonzero closed vector subgroups. Recalling that the connected component $c\left(H\right)$ of $H$ is equal to $H_{\mathbb{R}}\oplus H_{\mathbb{S}^{1}}$, we have that in this case $c\left(H\right)=H_{\mathbb{S}^{1}}$ is the connected component of zero.
Observe that $\pi^{-1}\left(H_{\mathbb{S}^{1}}\right)\subseteq G$ is a closed subgroup of $G=V\oplus A$ containing $V$. Therefore, we can write $\pi^{-1}\left(H_{\mathbb{S}^{1}}\right)=V\oplus B$ for some closed subspace $B$ of $A$. We have that $\pi|_{V\oplus B}:V\oplus B\rightarrow H_{\mathbb{S}^{1}}$ is surjective. Therefore, by Lemma 4.21 we must have that $H_{\mathbb{S}^{1}}$ is a torus group. By injectivity of $H_{\mathbb{S}^{1}}$, we have that $H\cong H_{\mathbb{S}^{1}}\oplus L$ for some non-Archimedean closed subgroup $L$ of $H$. Thus, if $C:=\pi^{-1}\left(L\right)$, then $C$ is a closed subgroup of $G$ containing $N$ such that $L\cong C/N$. Since non-Archimedean Polish groups form a thick subcategory of the category of pro-Lie Polish abelian groups, this implies that $L$ is non-Archimedean, and hence $L\subseteq A$. Furthermore, the composition
$A\overset{\pi|_{A}}{\rightarrow}H_{\mathbb{S}^{1}}\oplus L\rightarrow L\text{,}$
where $H_{\mathbb{S}^{1}}\oplus L\rightarrow L$ is the second coordinate projection, is a surjective continuous homomorphism. Since $A$ in injective in $\mathbf{PAb}_{\mathrm{nA}}$, the same holds for $L$, considering that $\mathbf{PAb}_{\mathrm{nA}}$ has enough injective objects and homological dimension one. This concludes the proof. ∎
@@
Without loss of generality, we can assume that $C$ has no nonzero closed vector subgroups. By [21, Proposition 3.8], there exists a locally compact Polish group $D$ containing $C$ as an open subgroup such that $D/C$ is a countable torsion group. Consider the pushout diagram
@@
For every locally compact Polish group $C$ there exists a continuous surjective homomorphism $G\rightarrow C$ for some element $C$ of $\mathcal{C}$.
@@
Without loss of generality, we can assume that $C$ has no vector groups as a direct summands, whence $C/C_{\mathbb{S}^{1}}$ is totally disconnected. We have that $C_{\mathbb{S}^{1}}^{\vee}$ is a countable torsion-free group. Thus there exists a short exact sequence $E\rightarrow C_{\mathbb{S}^{1}}^{\vee}\rightarrow S$ where $E$ is a countable free abelian group and $S$ is a countable torsion group. By duality, this gives a short exact sequence $A\rightarrow C_{\mathbb{S}^{1}}\rightarrow T$, where $A=S^{\vee}$ is a profinite abelian Polish group and $T=E^{\vee}$ is a torus group. Considering the inclusion $C_{\mathbb{S}^{1}}\rightarrow C$ one obtains by pushout a diagram
@@
By injectivity of $V$, we can write $C^{\prime}=V\oplus C^{\prime\prime}$, where we have an extension $A\rightarrow C^{\prime\prime}\rightarrow C/C_{\mathbb{S}^{1}}$. Hence, $C^{\prime\prime}$ is totally disconnected. The conclusion thus follows from Corollary 4.24. ∎
@@
We can write $D=T\oplus V\oplus A$ where $T$ is a torus, $V$ is a vector group, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We can also write $C=W\oplus B$ where $W$ is a vector group and $B$ is totally disconnected. By injectivity of $T$, $V$, and $W$ in the category of pro-Lie Polish abelian groups, we have that $\mathrm{Ext}\left(C,D\right)\cong\mathrm{Ext}\left(B,A\right)$. By Theorem 4.15, we have that $A\cong\prod_{n\in\omega}A_{n}$ where $A_{n}$ is a countable divisible group. Thus,
@@
It follows from Proposition 2.8 in the case when $\mathcal{A}=\mathbf{LCPAb}$, $\mathcal{B}=\mathbf{proPAb}$, and $\mathcal{C}$ and $\mathcal{D}$ are the classes defined above, Proposition 4.26, Lemma 4.25, and Lemma 4.22, that $\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{K}^{b}\left(\mathbf{proPAb}\right)\rightarrow\mathrm{K}^{b}\left(\mathbf{PAb}\right)$ has a total right derived functor $\mathrm{RHom}:\mathrm{D}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{D}^{b}\left(\mathbf{proPAb}\right)\rightarrow\mathrm{D}^{b}\left(\mathbf{PAb}\right)$, and $\mathrm{H}^{0}\circ\mathrm{RHom:D}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{D}^{b}\left(\mathbf{proPAb}\right)\rightarrow\mathrm{LH}\left(\mathbf{PAb}\right)$ is a cohomological derived functor of $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{K}^{b}\left(\mathbf{proPAb}\right)\rightarrow\mathrm{LH}\left(\mathbf{PAb}\right)$.

### v2

@@ 4.3. Injective non-Archimedean Polish abelian groups
(2$\Rightarrow$(3) If $A$ is divisible, then $A\cong\mathrm{\mathrm{lim}}_{k}A_{k}$ where, for every $k\in\omega$, $A_{k}$ is countable, divisible, and torsion-free and $A_{k+1}\rightarrow A_{k}$ is surjective with divisible kernel (notice that the kernel of $A_{k+1}\rightarrow A_{k}$ is a pure subgroup of $A_{k+1}$ since $A_{k}$ is torsion-free, and since $A_{k+1}$ is torsion-free divisible, the same holds for the kernel of $A_{k+1}\rightarrow A_{k}$). Since countable divisible groups are injective for countable abelian groups, it follows that $A$ is isomorphic to the product of countable divisible groups.
@@
    The category $\mathbf{PAb}_{\mathrm{nA}}$ has enough injective objects and enough projective objects, and homological dimension $1$;
@@
(1) The same proof as Lemma 3.44 shows that, for every non-Archimedean Polish abelian group $A$ there exists a surjective continuous homomorphism $\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}\rightarrow A$. By Theorem 3.52 this implies that $\mathbf{PAb}_{\mathrm{nA}}$ has enough projective objects and homological dimension $1$. The same proof as Corollary 4.6 shows that $\mathbf{PAb}_{\mathrm{nA}}$ has enough injective objects.
(2) We have that $\mathbf{proLiePAb}_{\mathbb{Z}}$ has enough projectives by Lemma 3.43 and Theorem 4.13. Furthermore, it has homological dimension $1$ by Lemma 3.45. Suppose that $A$ is a type $\mathbb{Z}$ Polish abelian group. Suppose that $A\rightarrow D\rightarrow B$ is a short exact sequence where $D$ and $B$ are type $\mathbb{Z}$ Polish abelian groups, and $D$ is injective in $\mathbf{proLiePAb}_{\mathbb{Z}}$. By Theorem 4.16, we have that $D$ is divisible. Thus, for $p\in\mathbb{Z}$ we have that $\mathrm{Hom}\left(\mathbb{Z}/p\mathbb{Z},B\right)=0$ and $\mathrm{Ext}\left(\mathbb{Z}/p\mathbb{Z},D\right)=0$. This implies that $\mathrm{Ext}\left(\mathbb{Z}/p\mathbb{Z},A\right)=0$ and $A$ is divisible. ∎
@@ 4.4. Ext as a group with a Polish cover
By Theorem 3.57, a pro-Lie Polish abelian group is essentially injective if and only if $G$ is isomorphic to $T\oplus V\oplus A$ where $T$ is a torus, $V$ is a vector group, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We let $\mathcal{D}$ be the collection of essentially injective pro-Lie Polish abelian groups. By Lemma 3.56 and Theorem 4.15, every pro-Lie Polish abelian group is isomorphic to a closed subgroup of an element of $\mathcal{D}$. We will now show that the class $\mathcal{D}$ is closed under taking quotients by closed subgroups. Towards this goal, we isolate a few lemmas concerning $\mathrm{Ext}$ of countable abelian groups, regarded as a group with a Polish cover.
@@
Let $H$ be a pro-Lie Polish group. Then we have that the connected component $c\left(H\right)$ of zero in $H$ is $H_{\mathbb{S}^{1}}\oplus H_{\mathbb{R}}$, and $H/c\left(H\right)$ is non-Archimedean. Thus, we have that $H=V\oplus A$ for some vector group $A$ and non-Archimedean abelian Polish group $A$ if and only if $H_{\mathbb{S}^{1}}=0$.
Suppose that $G$ is an essentially injective pro-Lie Polish abelian group and $N$ is a closed subgroup of $G$. Set $H:=G/N$. Since $G$ is essentially injective, we have that $G=T\oplus V\oplus A$ where $V$ is a vector group, $T$ is a torus, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We need to prove that $G/N$ is essentially injective.
We have that $T/\left(T\cap N\right)$ is injective, being the quotient of a torus group by a closed subgroup. Furthermore, letting $\pi:G\rightarrow V\oplus A$ be the canonical quotient map, we have that $\pi^{-1}\left(\pi\left(N\right)\right)=N+T$ is closed in $G$, and hence $\pi\left(N\right)$ is closed is $V\oplus A$. Thus, we have a short exact sequence
@@
Since $N$ is a closed subgroup of $G=V\oplus A$, we have that $c\left(N\right)\subseteq V=c(G)$ and hence $c\left(N\right)=N_{\mathbb{R}}$ is a vector group and $N_{\mathbb{S}^{1}}=0$. Thus, we can write $N=N_{\mathbb{R}}\oplus\Xi$ where $\Xi$ is non-Archimedean. Since $N_{\mathbb{R}}\subseteq V$ is a vector group, we can assume without loss of generality, possibly after replacing $G$ with $G/N_{\mathbb{R}}$, that $N_{\mathbb{R}}=0$ and $N=\Xi$ is non-Archimedean.
Then we have that $N\cap V$ is a closed subgroup of $V$ and $V/\left(N\cap V\right)$ is injective in $\mathbf{PAb}$. Thus, $V/\left(N\cap V\right)=W\oplus T$ where $T$ is a torus group and $W$ is a vector group. Since $W$ is projective in $\mathbf{PAb}$, we can assume without loss of generality that $W=0$ and $V/\left(N\cap V\right)$ is a torus group.
Thus, we have that $N/\left(N\cap V\right)$ is a closed subgroup of $G/N$. Being also injective in $\mathbf{PAb}$, it is a direct summand. Thus, after replacing $G$ with $G/V$ and $N$ with $N/\left(N\cap V\right)$, we can assume that $G=A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. In this case, we have that also $G/N$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$, concluding the proof. ∎
@@
Without loss of generality, we can assume that $C$ has no nonzero closed vector subgroups. By [22, Proposition 3.8], there exists a locally compact Polish group $D$ containing $C$ as an open subgroup such that $D/C$ is a countable torsion group. Consider the pushout diagram
@@
For every locally compact Polish group $C$ there exists a continuous surjective homomorphism $G\rightarrow C$ for some element $G$ of $\mathcal{C}$.
@@
Without loss of generality, we can assume that $C$ has no nonzero closed vector groups as a direct summand, whence $C/C_{\mathbb{S}^{1}}$ is totally disconnected. We have that $C_{\mathbb{S}^{1}}^{\vee}$ is a countable torsion-free group. Thus there exists a short exact sequence $E\rightarrow C_{\mathbb{S}^{1}}^{\vee}\rightarrow S$ where $E$ is a countable free abelian group and $S$ is a countable torsion group. By duality, this gives a short exact sequence $A\rightarrow C_{\mathbb{S}^{1}}\rightarrow T$, where $A=S^{\vee}$ is a profinite abelian Polish group and $T=E^{\vee}$ is a torus group. Considering the inclusion $C_{\mathbb{S}^{1}}\rightarrow C$ one obtains by pushout a diagram
@@
By projectivity of $V$, we can write $C^{\prime}=V\oplus C^{\prime\prime}$, where we have an extension $A\rightarrow C^{\prime\prime}\rightarrow C/C_{\mathbb{S}^{1}}$. Hence, $C^{\prime\prime}$ is totally disconnected. The conclusion thus follows from Corollary 4.23. ∎
@@
We can write $D=T\oplus V\oplus A$ where $T$ is a torus, $V$ is a vector group, and $A$ is non-Archimedean and injective in $\mathbf{PAb}_{\mathrm{nA}}$. We can also write $C=W\oplus B$ where $W$ is a vector group and $B$ is totally disconnected. By injectivity of $T$ and $V$, and projectivity of $W$, in the category of pro-Lie Polish abelian groups, we have that $\mathrm{Ext}\left(C,D\right)\cong\mathrm{Ext}\left(B,A\right)$. By Theorem 4.15, we have that $A\cong\prod_{n\in\omega}A_{n}$ where $A_{n}$ is a countable divisible group. Thus,
@@
It follows from Proposition 2.12 in the case when $\mathcal{A}=\mathbf{LCPAb}$, $\mathcal{B}=\mathbf{proLiePAb}$, and $\mathcal{C}$ and $\mathcal{D}$ are the classes defined above, Proposition 4.25, Lemma 4.24, and Lemma 4.21, that $\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{K}^{b}\left(\mathbf{proLiePAb}\right)\rightarrow\mathrm{K}^{b}\left(\mathbf{PAb}\right)$ has a total right derived functor $\mathrm{RHom}:\mathrm{D}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{D}^{b}\left(\mathbf{proLiePAb}\right)\rightarrow\mathrm{D}^{b}\left(\mathbf{PAb}\right)$, and $\mathrm{H}^{0}\circ\mathrm{RHom:D}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{D}^{b}\left(\mathbf{proLiePAb}\right)\rightarrow\mathrm{LH}\left(\mathbf{PAb}\right)$ is a cohomological derived functor of $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathbf{LCPAb}\right)\times\mathrm{K}^{b}\left(\mathbf{proLiePAb}\right)\rightarrow\mathrm{LH}\left(\mathbf{PAb}\right)$.